OpenStudy (anonymous):

A bag contains 6 red marbles, 4 blue marbles, 7 green marbles, and 3 yellow marbles. A marble is drawn from the bag and is not replaced. Then a second marble is drawn. What is the probability that both marbles drawn were yellow?

OpenStudy (anonymous):

9/400 9/380 3/200 3/190

OpenStudy (johnweldon1993):

Well first...there were 20 marbles in the beginning...and 1 was chosen..so that probability of it being yellow..was 3 yellow marbles -------------- = 3/20 chance 20 total marbles Now...it wasn't replaced so there are only 19 marbles left..we still want a yellow...so the chance now is 3 yellow marbles --------------- = 3/19 chance 19 total marbles To find the chance of those both happening...you multiply the 2 probabilities... 3 3 9 -- x --- = ---- 20 19 380

OpenStudy (anonymous):

Thank you!!!

OpenStudy (johnweldon1993):

Anytime! :)

OpenStudy (anonymous):

no answer is D C(3,2) / C(20,2) = 3/190